OpenStudy (anonymous):

Decompose this rational expression into partial fractions. -6x/(x-6)(x+3)

OpenStudy (anonymous):

what have u tried so far?

OpenStudy (anonymous):

-3/(x-6)+-1/(x+3)

OpenStudy (anonymous):

\[\large \frac{-6x}{(x-6)(x+3)}\]

OpenStudy (anonymous):

so how did u get to that (your answer there)?

OpenStudy (anonymous):

it was the mutiple choise answer like the textbook iam using for my online class is confusing

OpenStudy (anonymous):

Just some help: What you wanna do is something like this,\[\frac{-6x}{(x-6)(x+3)} = \frac{A}{(x-6)}+\frac{B}{(x+3)}\]

OpenStudy (anonymous):

x-2/x+4=x=6/x-4

OpenStudy (anonymous):

and then solve for A and B

OpenStudy (anonymous):

ah thanks

OpenStudy (anonymous):

lemme know if you still have problems

OpenStudy (anonymous):

for this one qustion x-2/x+4=x-6/x-4 i got x=8

OpenStudy (anonymous):

question

OpenStudy (anonymous):

i get x = 8 as well

OpenStudy (anonymous):

it fits in with the equations

OpenStudy (anonymous):

x-6/x<-1

OpenStudy (anonymous):

x≤ -3 or x≥0

OpenStudy (anonymous):

right